# An insight on some properties of high order nonstandard linear multistep methods

### v1

@@ 5.2 The SEIR system
Table 4: The errors and the corresponding orders of the different multistep methods.
For the last test concerning the SEIR model, we compare the behavior of the nonstandard methods and the standard counterparts for two different timesteps (the unmentioned parameters remain the same) for $t\in[0,10]$. Here we used Runge-Kutta methods again with an appropriate order to approximate the initial values of the scheme: standard ones for the standard multistep methods and nonstandard ones (with appropriate choices of $\varphi$) for the nonstandard versions. In all three cases, we only plot the numerical solutions of $I^{n}$, but the others also show signs of similar behavior. In Figure 6, we can see that the standard methods behave in an unstable way: they become negative, while the nonstandard ones remain in the upper half plane for the whole process, even for larger timesteps. Also, for a small timestep, the solutions are very close to each other and both of them behave as expected.
Figure 6: The comparison of the standard and nonstandard methods. Upper left: SSPMS(4,2) and NSSPMS(4,2) with $\Delta t_{1}=10/12$ and $\Delta t_{2}=0.02$. Upper right: SSPMS(4,3) and NSSPMS(4,3) with $\Delta t_{1}=10/15$ and $\Delta t_{2}=0.02$. Lower: SSPMS(6,4) and NSSPMS(6,4) with $\Delta t_{1}=10/15$ and $\Delta t_{2}=0.02$. The solutions for $\Delta t_{2}$ time-step are very close to each other in all three cases.
As a final remark, we mention that because of the form of the numerical schemes, the sum $S^{n}+E^{n}+I^{n}+R^{n}$ is preserved for every $\Delta t>0$ for the standard and nonstandard methods alike, so such numerical tests are omitted from the paper.
6 Conclusions and discussion
In this work, we considered nonstandard multistep methods. By using a nonstandard version of the classical Taylor series (Lemma 1), we were able to show that (under some conditions for the denominator function $\varphi$) the nonstandard versions of these methods have the same order as their standard counterparts. Note that Lemma 1 might also be useful in the observation of other, high-order nonstandard methods that are constructed in a similar fashion, i.e., involving a generalized denominator function.
In Section 3.2, it was shown that (under some further conditions for function $\varphi$) the nonstandard methods preserve the boundedness (and as a special case, positivity) properties of the original, continuous problem. To achieve this goal, a bound for the timestep should be calculated under which the Euler method preserves these properties. We note that the elementary stable property was not considered, but theorems similar to those in [49] might be proven in the future for nonstandard multistep methods too. These results would involve the observation of the elementary stability property of standard multistep methods, which might turn out to be harder (in general, these are not even true [50], as mentioned in Section 2) - this could also be a topic for further research. On the other hand, the observation of forward Euler methods might turn out to be easier, and for some systems, the preservation of positivity ensures elementary stability (see, for example, the case of an integro-differential equation [58] and its delayed version [59]).
In Section 5, we performed some tests for two test problems and observed that the nonstandard linear multistep methods behave as expected. In further articles, the same methods could also be applied to more robust ordinary systems and also to partial differential equations. One could also compare the performance of these multistep methods with other, higher-order methods (for example, the ones mentioned in Section 1).
Another interesting fact was the high order of convergence in the case of the method NSSPMS(4,3) in Section 5.1, which might have been caused by the use of the exact solutions for the initial values of the scheme. A further investigation in this regard (e.g., the increase of the rate of convergence in the case of higher order approximations of the initial values) might also prove to be fruitful. Finally, it would also be beneficial to observe the conditions under which some nonstandard multistep methods are A-stable and check whether the barriers of Dahlquist (see e.g. [51]) also hold in these nonstandard cases.
Acknowledgement
The author would like to thank Prof. István Faragó for his insightful comments during the preparation of the article. This research has been supported by the National Research, Development and Innovation Office – NKFIH, grant no. K137699. The research reported in this paper is also part of project no. BME-NVA-02, implemented with the support provided by the Ministry of Innovation and Technology of Hungary from the National Research, Development and Innovation Fund, financed under the TKP2021 funding scheme.
Declarations
Conflict of interest The author declares that they have no conflict of interest.
Appendix A The coefficients of the different methods
  • Explicit SSPMS(4,2) [60], with $\mathcal{C}=\frac{2}{3}$:
  • Explicit SSPMS(4,3) [60], with $\mathcal{C}=\frac{1}{3}$:
  • Explicit SSPMS(6,4) [61, 62], with $\mathcal{C}=0.1648$:
[…]
  • Explicit SSPRK(2,2) with $\mathcal{C}=1$ [63]:
    $u^{(1)}$ $=u^{n}+\Delta tf(u^{n}),$
  • Explicit SSPRK(3,3) with $\mathcal{C}=1$ [52]:
    $u^{(1)}$ $=u^{n}+\Delta tf(u^{n}),$
    $u^{(2)}$ $=\dfrac{3}{4}u^{n}+\dfrac{1}{4}u^{(1)}+\dfrac{1}{4}\Delta tf(u^{(1)}),$
  • Explicit SSPRK(4,10) with $\mathcal{C}=6$ [62]:
    $u^{(1)}$ $=u^{n}+\dfrac{1}{6}\Delta tf(u^{n}),$
    $u^{(j+1)}$ $=u^{(j)}+\dfrac{1}{6}\Delta tf(u^{(j)}),\qquad j=1,2,3,$
    $u^{(5)}$ $=\dfrac{3}{5}u^{n}+\dfrac{2}{5}u^{(4)}+\dfrac{1}{15}\Delta tf(u^{(4)}),$
    $u^{(j+1)}$ $=u^{(j)}+\dfrac{1}{6}\Delta tf(u^{(j)}),\qquad j=5,\dots,8,$

### v2

@@ Appendix B. Tables of errors and orders
Table 5: The errors and the corresponding orders of the different multistep methods while solving the logistic equation with $c=500$, plotted in Figure 6.
$0.1$ $5.3467\cdot 10^{-1}$ - $5.1861\cdot 10^{-1}$ - $6.1021\cdot 10^{-1}$ -
$0.1\cdot 2^{-1}$ $3.2527\cdot 10^{-1}$ $0.7170$ $2.6609\cdot 10^{-2}$ $0.9627$ $4.5905\cdot 10^{-1}$ $0.4106$
$0.1\cdot 2^{-2}$ $1.5828\cdot 10^{-1}$ $1.0391$ $1823\cdot 10^{-1} 1.1$7031.1703 $2.7619\cdot 10^{-1}$ $0.7330$
$0.1\cdot 2^{-3}$ $7.2915\cdot 10^{-2}$ $1.1181$ $5.3255\cdot 10^{-2}$ $1.1506$ $1.4172\cdot 10^{-1}$ $0.9625$
$0.1\cdot 2^{-4}$ $3.4220\cdot 10^{-2}$ $1.0913$ $2.4977\cdot 10^{-2}$ $1.0923$ $6.8670\cdot 10^{-2}$ $1.0453$
$0.1\cdot 2^{-5}$ $1.6476\cdot 10^{-2}$ $1.0544$ $1.2060\cdot 10^{-2}$ $1.0503$ $3.3198\cdot 10^{-2}$ $1.0486$
$0.1\cdot 2^{-6}$ $8.0718\cdot 10^{-3}$ $1.0294$ $5.9220\cdot 10^{-3}$ $1.0262$ $1.6230\cdot 10^{-2}$ $1.0324$
$0.1\cdot 2^{-7}$ $3.9934\cdot 10^{-3}$ $1.0153$ $2.9338\cdot 10^{-3}$ $1.0133$ $8.0116\cdot 10^{-3}$ $1.0185$
$0.1\cdot 2^{-8}$ $1.9859\cdot 10^{-3}$ $1.0078$ $1.4601\cdot 10^{-3}$ $1.0067$ $3.9785\cdot 10^{-3}$ $1.0099$
$0.1\cdot 2^{-9}$ $9.9029\cdot 10^{-4}$ $1.0039$ $7.2833\cdot 10^{-4}$ $1.0034$ $1.9823\cdot 10^{-3}$ $1.0051$
$0.1$ $5.6656\cdot 10^{-1}$ - $5.1799\cdot 10^{-1}$ - $5.3549\cdot 10^{-1}$ -
$0.1\cdot 2^{-1}$ $3.4930\cdot 10^{-1}$ $0.6977$ $2.4090\cdot 10^{-1}$ $1.1044$ $2.8795\cdot 10^{-1}$ $0.8950$
$0.1\cdot 2^{-2}$ $1.3496\cdot 10^{-1}$ $1.3720$ $6.5892\cdot 10^{-2}$ $1.8703$ $9.2618\cdot 10^{-2}$ $1.6364$
$0.1\cdot 2^{-3}$ $3.8084\cdot 10^{-2}$ $1.8253$ $1.6146\cdot 10^{-2}$ $2.0289$ $2.3895\cdot 10^{-2}$ $1.9546$
$0.1\cdot 2^{-4}$ $9.7659\cdot 10^{-3}$ $1.9634$ $4.0000\cdot 10^{-3}$ $2.0131$ $5.9821\cdot 10^{-3}$ $1.9980$
$0.1\cdot 2^{-5}$ $2.4566\cdot 10^{-3}$ $1.9911$ $9.9818\cdot 10^{-4}$ $2.0026$ $1.4962\cdot 10^{-3}$ $1.9993$
$0.1\cdot 2^{-6}$ $6.1532\cdot 10^{-4}$ $1.9972$ $2.4954\cdot 10^{-4}$ $2.0000$ $3.7424\cdot 10^{-4}$ $1.9992$
$0.1\cdot 2^{-7}$ $1.5393\cdot 10^{-4}$ $1.9990$ $6.2399\cdot 10^{-5}$ $1.9997$ $9.3594\cdot 10^{-5}$ $1.9994$
$0.1\cdot 2^{-8}$ $3.8495\cdot 10^{-5}$ $1.9996$ $1.5602\cdot 10^{-5}$ $1.9998$ $2.3403\cdot 10^{-5}$ $1.9997$
$0.1\cdot 2^{-9}$ $9.6252\cdot 10^{-6}$ $1.9998$ $3.9010\cdot 10^{-6}$ $1.9998$ $5.8515\cdot 10^{-6}$ $1.9998$
$0.1$ $5.2069\cdot 10^{-1}$ - $5.1734\cdot 10^{-1}$ -
$0.1\cdot 2^{-1}$ $2.3401\cdot 10^{-1}$ $1.1539$ $2.1140\cdot 10^{-1}$ $1.2911$
$0.1\cdot 2^{-2}$ $4.4870\cdot 10^{-2}$ $2.9092$ $2.5323\cdot 10^{-2}$ $3.8940$
$0.1\cdot 2^{-3}$ $5.9732\cdot 10^{-3}$ $2.9091$ $1.7033\cdot 10^{-3}$ $3.8940$
$0.1\cdot 2^{-4}$ $7.5410\cdot 10^{-4}$ $2.9857$ $1.0731\cdot 10^{-4}$ $3.9885$
$0.1\cdot 2^{-5}$ $9.4509\cdot 10^{-5}$ $2.9962$ $6.7211\cdot 10^{-6}$ $3.9969$
$0.1\cdot 2^{-6}$ $1.1826\cdot 10^{-5}$ $2.9985$ $4.2047\cdot 10^{-7}$ $3.9986$
$0.1\cdot 2^{-7}$ $1.4789\cdot 10^{-6}$ $2.9993$ $2.6294\cdot 10^{-8}$ $3.9992$
$0.1\cdot 2^{-8}$ $1.8491\cdot 10^{-7}$ $2.9996$ $1.6500\cdot 10^{-9}$ $3.9942$
$0.1\cdot 2^{-9}$ $2.3129\cdot 10^{-8}$ $2.9991$ $1.1539\cdot 10^{-10}$ $3.8379$
Table 6: The errors and the corresponding orders of method NSSPMS(6,4) using different $\varphi$ functions applied to the SEIR model with $\Pi=0$, plotted in Figure 9.
NSSPMS(4,2) NSSPMS(4,3) with $\varphi_{8}$ NSSPMS(4,3) with $\varphi_{7}$
$\Delta t$ errors orders errors orders errors orders
$0.05$ $2.4440\cdot 10^{-3}$ - $2.0610\cdot 10^{-2}$ - $3.4765\cdot 10^{-2}$ -
$0.05\cdot 2^{-1}$ $2.6209\cdot 10^{-4}$ $3.2210$ $1.5549\cdot 10^{-3}$ $3.7284$ $5.4335\cdot 10^{-3}$ $2.6776$
$0.05\cdot 2^{-2}$ $3.6849\cdot 10^{-5}$ $2.8303$ $9.9708\cdot 10^{-5}$ $3.9630$ $7.1272\cdot 10^{-4}$ $2.9305$
$0.05\cdot 2^{-3}$ $6.9473\cdot 10^{-6}$ $2.4071$ $6.2038\cdot 10^{-6}$ $4.0065$ $9.0404\cdot 10^{-5}$ $2.9789$
$0.05\cdot 2^{-4}$ $1.7148\cdot 10^{-6}$ $2.0184$ $3.7664\cdot 10^{-7}$ $4.0419$ $1.1370\cdot 10^{-5}$ $2.9912$
$0.05\cdot 2^{-5}$ $4.2660\cdot 10^{-7}$ $2.0071$ $2.1926\cdot 10^{-8}$ $4.1025$ $1.4253\cdot 10^{-6}$ $2.9958$
$0.05\cdot 2^{-6}$ $1.0643\cdot 10^{-7}$ $2.0030$ $1.1617\cdot 10^{-9}$ $4.2384$ $1.7842\cdot 10^{-7}$ $2.9979$
$0.05\cdot 2^{-7}$ $2.6581\cdot 10^{-8}$ $2.0014$ $5.0217\cdot 10^{-11}$ $4.5318$ $2.2318\cdot 10^{-8}$ $2.9990$
$0.05\cdot 2^{-8}$ $6.6424\cdot 10^{-9}$ $2.0006$ $2.9863\cdot 10^{-12}$ $4.0717$ $2.7908\cdot 10^{-9}$ $2.9995$
NSSPRK(2,2) NSSPRK(3,3) with $\varphi_{8}$ NSSPRK(3,3) with $\varphi_{7}$
$\Delta t$ errors orders errors orders errors orders
$0.05$ $4.2992\cdot 10^{-4}$ - $3.1523\cdot 10^{-4}$ - $1.7140\cdot 10^{-3}$ -
$0.05\cdot 2^{-1}$ $4.8459\cdot 10^{-5}$ $3.1492$ $1.9170\cdot 10^{-5}$ $4.0395$ $2.1592\cdot 10^{-4}$ $2.9888$
$0.05\cdot 2^{-2}$ $1.1345\cdot 10^{-5}$ $2.0947$ $1.1269\cdot 10^{-6}$ $4.0885$ $2.7019\cdot 10^{-5}$ $2.9984$
$0.05\cdot 2^{-3}$ $2.7744\cdot 10^{-6}$ $2.0318$ $6.1496\cdot 10^{-8}$ $4.1957$ $3.3780\cdot 10^{-6}$ $2.9997$
$0.05\cdot 2^{-4}$ $6.8803\cdot 10^{-6}$ $2.0116$ $2.7269\cdot 10^{-9}$ $4.4952$ $4.2228\cdot 10^{-7}$ $2.9999$
$0.05\cdot 2^{-5}$ $1.7145\cdot 10^{-7}$ $2.0047$ $1.9683\cdot 10^{-10}$ $3.7923$ $5.2786\cdot 10^{-8}$ $3.0000$
$0.05\cdot 2^{-6}$ $4.2800\cdot 10^{-8}$ $2.0021$ $2.7238\cdot 10^{-11}$ $2.8532$ $6.5982\cdot 10^{-9}$ $3.0000$
$0.05\cdot 2^{-7}$ $1.0693\cdot 10^{-8}$ $2.0010$ $3.4914\cdot 10^{-12}$ $2.9637$ $8.2477\cdot 10^{-10}$ $3.0000$
$0.05\cdot 2^{-8}$ $2.6723\cdot 10^{-9}$ $2.0005$ $5.0260\cdot 10^{-13}$ $2.7963$ $1.0307\cdot 10^{-10}$ $3.0000$
NSSPRK(6,4) NSSPRK(10,4)
$\Delta t$ errors orders errors orders
$0.05$ $1.1739\cdot 10^{-1}$ - $2.4392\cdot 10^{-7}$ -
$0.05\cdot 2^{-1}$ $2.1800\cdot 10^{-2}$ $2.4289$ $1.5246\cdot 10^{-8}$ $3.9999$
$0.05\cdot 2^{-2}$ $1.6444\cdot 10^{-3}$ $3.7287$ $9.5298\cdot 10^{-10}$ $3.9999$
$0.05\cdot 2^{-3}$ $1.0584\cdot 10^{-4}$ $3.9576$ $5.9574\cdot 10^{-11}$ $3.9997$
$0.05\cdot 2^{-4}$ $6.6819\cdot 10^{-6}$ $3.9855$ $3.7333\cdot 10^{-12}$ $3.9962$
$0.05\cdot 2^{-5}$ $4.1960\cdot 10^{-7}$ $3.9932$ $2.4303\cdot 10^{-13}$ $3.9413$
$0.05\cdot 2^{-6}$ $2.6286\cdot 10^{-8}$ $3.9966$ $5.3291\cdot 10^{-14}$ $2.1892$
$0.05\cdot 2^{-7}$ $1.6446\cdot 10^{-9}$ $3.9985$ $6.7391\cdot 10^{-14}$ -
$0.05\cdot 2^{-8}$ $1.0248\cdot 10^{-10}$ $4.0043$ $1.0170\cdot 10^{-13}$ -
Table 7: The errors and the corresponding orders of the different nonstandard methods applied to the SEIR model with $\Pi=0$, plotted in Figure 10.